Exercise 2-15 (Algo) Plantwide and Departmental Predetermined Overhead Rates; Job Costs [LO2-1, LO2-2, LO2-3, LO2-4]

[The following information applies to the questions displayed below.]

Delph Company uses a job-order costing system and has two manufacturing departments—Molding and Fabrication. The company provided the following estimates at the beginning of the year:

  

Molding Fabrication Total
Machine-hours 24,000 32,000 56,000
Fixed manufacturing overhead cost $ 740,000 $ 200,000 $ 940,000
Variable manufacturing overhead cost per machine-hour $ 3.00 $ 1.50

  

During the year, the company had no beginning or ending inventories and it started, completed, and sold only two jobs—Job D-70 and Job C-200. It provided the following information related to those two jobs:

  

Job D-70 Molding Fabrication Total
Direct materials cost $ 370,000 $ 320,000 $ 690,000
Direct labor cost $ 240,000 $ 160,000 $ 400,000
Machine-hours 16,000 8,000 24,000

  

Job C-200 Molding Fabrication Total
Direct materials cost $ 220,000 $ 240,000 $ 460,000
Direct labor cost $ 140,000 $ 240,000 $ 380,000
Machine-hours 8,000 24,000 32,000

Delph had no underapplied or overapplied manufacturing overhead during the year.

rev: 07_21_2020_QC_CS-217627

Exercise 2-15 (Algo) Part 1

Required:

1. Assume Delph uses departmental predetermined overhead rates based on machine-hours.

a. Compute the departmental predetermined overhead rates.

b. Compute the total manufacturing cost assigned to Job D-70 and Job C-200.

c. If Delph establishes bid prices that are 150% of total manufacturing cost, what bid prices would it have established for Job D-70 and Job C-200?

d. What is Delph’s cost of goods sold for the year?
Answer #1

Requirement 1:

a. Departmental predetermined overhead rate

Departmental predetermined overhead rate = Total Departmental manufacturing overhead cost / Total departmental machine hours

Molding

Variable overhead cost = 24,000 machine hours * $3.00 per machine hour = $72,000

Fixed overhead cost = $740,000

Total Manufacturing overhead cost = $740,000 + $72,000 = $812,000

Departmental predetermined overhead rate = $812,000 / 24,000 machine hours = $33.83 per machine hour

Fabrication

Variable overhead cost = 32,000 machine hours * $1.50 per machine hour = $48,000

Fixed overhead cost = $200,000

Total Manufacturing overhead cost = $200,000 + $48,000 = $248,000

Departmental predetermined overhead rate = $248,000 / 32,000 machine hours = $7.75 per machine hour

b. Total Manufacturing cost assigned to Jobs

Job D-70

(i) Manufacturing overheads

Molding department's overhead cost = 16,000 machine hours * $33.83 per machine hour = $541,280

Fabrication department's overhead cost = 8,000 machine hours * $7.75 per machine hour = $62,000

Total manufacturing overhead cost assigned = $541,280 + $62,000 = $603,280

(ii) Total Manufacturing Cost

Particulars $
Direct Materials 690,000
Direct labor cost 400,000
Manufacturing overheads 603,280
Total Manufacturing Overhead cost 1,693,280

Job C-200

(i) Manufacturing overheads

Molding department's overhead cost = 8,000 machine hours * $33.83 per machine hour = $270,640

Fabrication department's overhead cost = 24,000 machine hours * $7.75 per machine hour = $186,000

Total manufacturing overhead cost = $270,640 + $186,000 = $456,640

(ii) Total Manufacturing Cost

Particulars $
Direct Materials 460,000
Direct labor cost 380,000
Manufacturing overheads 456,640
Total Manufacturing Overhead cost 1,296,640

c. Bid Price

Bid price = 150% of total manufacturing cost

Job D-70

Bid Price = $1,693,280 * 150% = $2,539,920

Job C-200

Bid Price = $1,296,640 * 150% = $1,944,960

d. Delph's cost of goods sold

Delph's cost of goods sold = $1,693,280 + $1,296,640 = $2,989,920

All the best...